Starting from a population of 100 individuals in which 100 A alleles and 100 a alleles are present, what level of heterozygosity would be most likely after 10,000 generations?

a.) 0.0
b.) 0.5
c.) 1.0
d.) 0.05
Answer #1

a). 0.0

The frequency of A allele is = 100/200 = 0.5

The frequency of a allele is = 100/200 = 0.5

The frequency of heterozygotes, is = 2*0.5*0.5 = 0.5

The heterozygosity loss after “t” number of generations can be calculated by the below formula.

Ho 0

Where, Ht = heterozygosity after t generations

H0 = Initial heterozygosity

t = number of generations

Ne = Effective population size.

Assume that the total number of population is = 100

Then the number of heterozygous individuals, H0 = 50

Ht = 50(1-1/2*100)^10000

Ht = 50 (1- 0.005)^100000

The heterozygosis lost after 10, 000 generations is 100%. Means, the heterozygosity is zero.